What concentration of NaCHO2 is required to produce a buffer with a pH of 4.25 in a solution that is 1.32M in HCHO2? Ka = 1.8 x 10-4
Answer #1

Ka = 1.8*10^-4

use:

pKa = -log Ka

= -log (1.8*10^-4)

= 3.74

Use:

pH = pKa + log ([CHO2-]/[HCHO2])

4.25 = 3.74 + log ([CHO2-]/1.32)

log ([CHO2-]/1.32) = 0.51

[CHO2-]/1.32 = 10^0.51

[CHO2-]/1.32 = 3.24

[CHO2-] = 4.27 M

Answer: 4.27 M